# Decoherent histories with(out) objectivity in a (broken) apparatus

### v1

@@ Abstract
The decoherence program [1, 2, 3, 4] is fundamentally motivated by a measurement problem: when describing measurement as a unitary system-apparatus interaction [5], their bipartite entanglement cannot determine the measurement basis. However, the apparatus (more generally, the environment) has many degrees of freedom, and the multipartite correlation may select a set of preferred “pointer states”, through the proliferation of classical records [6, 7]. In decoherence, and its refinement quantum Darwinism [8, 9, 10, 11, 12, 13, 14], classicality is defined as objectivity [15, 16, 17, 18], the redundant accessibility of records.
@@
Model. Let us start by one of the simplest unitary models of measurement [5], where both the apparatus $A$ and the system $S$ are a qubit. Initially, they are disentangled and the apparatus is always in a “ready” state $|0\rangle_{A}$. The measurement unitary is such that $|i\rangle_{S}|0\rangle_{A}\mapsto|i\rangle_{S}|i\rangle_{A}$, for $i=0,1$, so that measuring the superposition state $(|0\rangle_{S}+|1\rangle_{S})/\sqrt{2}$ results in a maximally entangled pair:
@@
From $|\Psi\rangle$, it is impossible to know in which basis the measurement took place [6, 7].
In our more realistic apparatus model [49, 50, 51], $|\Psi\rangle$ is the initial state of a cascade process involving more and more qubits, bridging micro- and macroscopic realms. Starting with the qubit $A$, at each time step, every existing qubit interacts with a new qubit in a factorized state, so that there are $N_{t}=2^{t}$ qubits after $t$ steps. (During all this, $S$ remains intact.) Such an inflationary dynamics is known as “concatenation” in the error correction code literature [52, 53], and often represented diagrammatically by a tree, see Fig. 1 [54, 55]. Every node corresponds to an isometry that outputs two qubits from one, $v:\mathbb{C}^{2}\to\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, such that $v^{\dagger}v=I$. The evolution from time step $t$ to $t+1$ is given by $V_{t+1,t}=v^{\otimes 2^{t}}$. We also denote $V_{t,s}:=V_{t,t-1}V_{t-1,t-2}\dots V_{s+1,s}$ for $t\geq s$. We will focus on the following concrete family of models:
@@
where $X$ is the Pauli-$x$ matrix, and $\theta\in(0,\pi/2)$ is the tuning parameter. When $\theta=0$, the model realizes a repetition code, which clearly has objectivity: every single apparatus qubit has perfect classical correlation with the system qubit. Increasing $\theta$ perturbs the code, and, as we shall see, eventually destroys the objectivity beyond some threshold
@@
Phase diagram. It is known [48] that one may locate $\theta_{c}$ by considering the correlation between $S$ and an extensive observable of the apparatus. Here we point out a general way to do this, in connection with the multi-scale entanglement renormalization Ansatz (MERA) [56, 57, 58, 59], of which the tree models are a special case. The Heisenberg evolution of operators is known to implement their renormalization group (RG) at a fixed point, whose RG data is encoded by the local tensors [58]. Here, the only local tensor is the isometry $v$. It defines the scaling operators $O_{x}$ and their scaling dimension $x$ as follows:
$v^{\dagger}(O_{x}\otimes\mathbf{1})v=v^{\dagger}(\mathbf{1}\otimes O_{x})v=2^{-x}O_{x}\,.$ (4)
(We embed the tree in $d=1$ spatial dimension.) $v$ also determines the operator product expansion (OPE), $v^{\dagger}(O_{x}\otimes O_{x^{\prime}})v=\sum_{x^{\prime\prime}}C_{xx^{\prime}}^{x^{\prime\prime}}O_{x^{\prime\prime}}\,.$ For the model (2), there are only two scaling operators with finite scaling dimension, the identity $I$ with $x=0$, and a nontrivial one:
$x=-\log_{2}(\cos\theta)$ (5)
with $O_{x}=Z\cos(\theta/2)+Y\sin(\theta/2).$ Its only OPE is to identity $C_{xx}^{0}=(\cos\theta)^{2}>0$.
We can then compute correlation functions by RG [58, 60]. For concreteness, consider the extensive observable:
$\mathcal{O}_{t}=\sum_{j=1}^{N_{t}}Z_{j}$ (6)
where $N_{t}=2^{t}$ and $Z_{j}$ acts on the $j$-th qubit at time $t$ (the results below apply to any local operator that contains $O_{x}$). Its correlation with any operator $O^{\prime}$ acting on $S$ has the following scaling behavior
$\mu_{t}:=\langle O^{\prime}_{S}V^{\dagger}_{t,0}\mathcal{O}_{t}\mathcal{V}_{t,0}\rangle\sim N_{t}^{1-x}$ (7)
where $\left<[\dots]\right>=\left<\Psi|[\dots]|\Psi\right$>. Meanwhile, the second moment of $\mathcal{O}$ scales as
$\eta^{2}_{t}:=$ $\langle V^{\dagger}_{t,0}\mathcal{O}_{t}^{2}V_{t,0}\rangle=\langle V^{\dagger}_{t,0}\sum_{ij}Z_{i}Z_{j}V_{t,0}\rangle$
[…]
There is a non-analytical change at the threshold $x_{c}=1/2$, which gives (3) via (5), because $\eta^{2}_{t}$ is dominated by pairs $Z_{i}Z_{j}$ separated by graph distance $\sim t$ when $x<1/2$, but $\sim 1$ when $x>1/2$ (see [61, 62, 63] for other “inference” problems where the threshold $x_{c}=d/2$ appears). It is not hard to show that the scaling laws (7) and (8) still hold if the sum in (6) is over any fixed fraction of the apparatus [48].
We determine the phase diagram by comparing the noise $\eta_{t}$ with the signal $\mu_{t}$ . When $x>1/2$ ($\theta>\theta_{c}$), the noise dominates as $t\gg 1$. This is the encoding phase where information on $S$ becomes inaccessible. When $x<1/2$, noise and signal are comparable. This is the apparatus phase where information on $S$ is accessible from a coarse-grained observation of the apparatus and of its fractions. Therefore objectivity emerges. We will further describe what information is inferred below.
@@
$|\vec{m}\rangle:=|K^{T}_{m_{T}}V_{T,T-1}\dots V_{\tau+1,\tau}K^{1}_{m_{\tau}}V_{\tau,0}|\Psi\rangle\,.$ (9)
@@
In the sense of (13), classicality à la decoherent histories amounts to non-disturbance by third-party measurements (which is also part of macro-realism [64]). Namely, classicality emerges if the existence of measurements in $t\notin\mathbf{t}$ do not affect the outcome distribution of the ones in $t\in\mathbf{t}$. This is a reasonable way to define “classical” deal measurements are non-invasive in classical physics, while any decent quantum mechanics textbook discusses the inference pattern being destroyed by monitoring the double slits.
@@
$K^{t}_{m}=\frac{1}{(2\pi\Gamma^{2})^{1/4}}e^{-(\mathcal{O}_{t}/\eta_{t}-m)^{2}/(4\Gamma^{2})}\,,m\in\mathbb{R},$ (14)
@@
Figure 2: Decoherent histories probe $\|\Delta^{\mathbf{t}}\|_{1}$ with $\mathbf{t}=\{L+\tau\}$, as function of $\tau$ (measurement starting time), for $\theta=0.15\pi<\theta_{c}$ (a) and $\theta=0.3\pi>\theta_{c}$ (b). Approximate decoherent histories emerge for coarse-grained measurements ($\Gamma=0.1$) and $\tau\gg 1$, but not for $\tau\sim 1$ or for fine-grained measurements (fg, $\Gamma\eta_{t}=0.1$). Insets (i, ii) show the pdf $p^{\mathbf{t}}$ (no third-party) and marginal distributions in non-decoherent scenarios ($T=11$). Straight lines show decay rate of $\|\Delta^{\mathbf{t}}\|_{1}$ (15), and (53) in [60]. (a) For $\theta<\theta_{c}$, $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{2\tau(x-1)}$ with no $L$ scaling. (b) For $\theta>\theta_{c}$, : $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{-T}\max(1,2\cos\theta)^{2L}$. Inset (iii): $L$-dependence is shown for two $\theta>\theta_{c}$ ($\tau=6$).
@@
Histories statistics. Although decoherent histories emerge at $t\geq\tau\gg 1$ independently of objectivity, they have distinct statistics $p_{\vec{m}}$ (10) in the two phases. In the encoding phase ($x>1/2$), the time sequence $(m_{t})$ is a Gaussian process with finite temporal correlation:
$\mathbb{E}[m_{t}]=0,\mathbb{E}[m_{t}m_{t^{\prime}}]\sim e^{-\kappa|t-t^{\prime}|}\,,\,\kappa=(x-1/2)\ln 2$ (17)
for $|t-t^{\prime}|\gg 1$. The correlation time $\kappa^{-1}$ diverges near the threshold $x=1/2$. Meanwhile, the apparatus phase ($x<1/2$) is characterized by non-ergodic histories with the following probabilistic law:
@@
where $\xi_{t}$ are i.i.d standard Gaussian and $\mathcal{M}$ is a random variable with the same distribution as $\mathcal{O}_{t}/\eta_{t}$ in the long time limit (this distribution is non-Gaussian, see Fig. 2):
$\forall f\,,\,\overline{f(\mathcal{M})}=\lim_{t\to\infty}\langle V_{t,0}^{\dagger}f(\mathcal{O}_{t}/\eta_{t})V_{t,0}\rangle.$ (19)
Namely, each history instance freezes around a random value, just like an apparatus pointer. A classical analogue of the two behaviors is an Ornstein-Uhlenbeck (OU) [69] process $\mathrm{d}M_{t}=-\kappa M_{t}\mathrm{d}t+\mathrm{d}W_{t}$ with $\kappa>0$ (encoding) or $\kappa<0$ (apparatus). In the latter case, $m_{t}=e^{\kappa t}M_{t}$ also freezes and is non-Gaussian if the initial condition $m_{t=0}$ is so.

### v2

@@ Abstract
The decoherence program [1, 2, 3, 4] is fundamentally motivated by a measurement problem, namely how to determine “what is being measured” when describing measurement as a unitary interaction between the system and the apparatus [5] . The key idea is that, the apparatus, and more generally, the environment) has many degrees of freedom, which allow to select a set of preferred “pointer states”, through the proliferation of classical records [6, 7]. In decoherence, and its refinement quantum Darwinism [8, 9, 10, 11, 12, 13, 14], classicality is defined as objectivity [15, 16, 17, 18], the redundant accessibility of records.
@@
Model and objectivity threshold. Consider one of the simplest unitary models of measurement [5], where both the apparatus $A$ and the system $S$ are a qubit. Initially, they are disentangled and the apparatus is always in a “ready” state $|0\rangle_{A}$. The measurement unitary is such that $|i\rangle_{S}|0\rangle_{A}\mapsto|i\rangle_{S}|i\rangle_{A}$, for $i=0,1$, so that measuring the superposition state $(|0\rangle_{S}+|1\rangle_{S})/\sqrt{2}$ results in a maximally entangled pair:
@@
This is a situation where the aforementioned measurement basis problem is the most acute: it is impossible to know in which basis the measurement took place [6, 7].
In our apparatus model, $|\Psi\rangle$ is the initial state of a cascade process involving more and more qubits, bridging micro- and macroscopic realms [49, 50, 51]. Starting with the qubit $A$, at each time step, every existing qubit interacts with a new qubit (in a pure state), so that there are $N_{t}=2^{t}$ qubits after $t$ steps. During all this, $S$ remains intact. Such an inflationary dynamics is known as “concatenation” in the error correction code literature [52, 53, 54], and often represented diagrammatically by a tree, see Fig. 1 [55, 56]. Every node corresponds to an isometry that outputs two qubits from one, $v:\mathbb{C}^{2}\to\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, such that $v^{\dagger}v=I$. The evolution from time step $t$ to $t+1$ is given by $V_{t+1,t}=v^{\otimes 2^{t}}$. We also denote $V_{t,s}:=V_{t,t-1}V_{t-1,t-2}\dots V_{s+1,s}$ for $t\geq s$. We will focus on the following concrete family of models:
@@
where $X$ is the Pauli-$x$ matrix, and $\theta\in(0,\pi/2)$ is the tuning parameter. Our model is equivalent to a quantum circuit with non-local control gates [57], similar to those realized by coherently transportable neutral atoms [58].
When $\theta=0$, our model realizes a repetition code, which clearly has objectivity: every single apparatus qubit has perfect classical correlation with the system qubit. Increasing $\theta$ perturbs the code, and, as we shall see, destroys the objectivity beyond some threshold $\theta_{c}$. To determine $\theta_{c}$, we consider the correlation between $S$ and an extensive observable of the apparatus, the total magnetization
$\mathcal{Q}_{t}=\sum_{j=1}^{N_{t}}Z_{j}.$ (3)
where $N_{t}=2^{t}$ and $Z_{j}$ acts on the $j$-th qubit at time $t$. At $\theta=0$, $\mathcal{Q}_{t}$ is also perfectly correlated with $Z_{S}$, the $Z$ operator on $S$. Exploiting the tree structure (see below and [57] for methods), we calculate the correlation between $O^{\prime}_{S}$ (for any $O^{\prime}$) and $\mathcal{Q}_{t}$ for any $\theta$. As $t\to\infty$, we find the following scaling behavior
$\mu_{t}:=\langle Z_{S}V^{\dagger}_{t,0}\mathcal{Q}_{t}V_{t,0}\rangle\sim N_{t}^{1-x_{1}}$ (4)
where $\left<[\dots]\right>=\left<\Psi|[\dots]|\Psi\right$>, and the exponent is
$x_{1}=-\log_{2}(\cos\theta).$ (5)
We also compute the scaling of the second moment of $\mathcal{Q}_{t}$:
$\eta^{2}_{t}:=$ $\langle V^{\dagger}_{t,0}\mathcal{Q}_{t}^{2}V_{t,0}\rangle=\langle V^{\dagger}_{t,0}\sum_{ij}Z_{i}Z_{j}V_{t,0}\rangle$
$\sim$ $\begin{cases}N_{t}&x_{1}>1/2\,,\\ N_{t}^{2(1-x_{1})}&x_{1}<1/2\,.\end{cases}$ (6)
The threshold corresponds to $x_{1}=1/2$ or
@@
When $\theta<\theta_{c}$, $x_{1}<1/2$, $\mu_{t}\sim\eta_{t}$, the correlation between $\mathcal{Q}_{t}$ and any operator on $S$ is comparable to its fluctuation amplitude. Moreover, (4) and (6) hold as well if $\mathcal{Q}_{t}$ (3) is replaced by the total magnetization of a fixed fraction $fN_{t}$ of apparatus qubits [57]. Thus, objectivity prevails: information about $S$ can be obtained from small fractions of the apparatus. We call $\theta<\theta_{c}$ the “apparatus phase”. When $\theta>\theta_{c},x_{1}>1/2$, the correlation is vanishing compared to the noise, $\mu_{t}\ll\eta_{t}$, and observing $\mathcal{Q}_{t}$ yields no information about the qubit. This is the “encoding phase”. Based on previous work on a very similar model [48], we expect that the Holevo quantity between an apparatus fraction and $S$ vanishes if $\theta>\theta_{c}$ as $t\to\infty$. In other words, information about $S$ is inaccessible to other observables as well, not just $\mathcal{Q}_{t}$.
We remark that our expanding tree model is a special case of the multi-scale entanglement renormalization Ansatz (MERA) tensor networks [59, 60, 61, 62], which are known to generate scale invariant states. In that context, the exponent $x_{1}$ above is the lightest scaling dimension. Interestingly, the threshold phenomena at $x_{1}=1/2$ appeared in various recently studied scale invariant “inference” problems [63, 64, 65].
@@
$|\vec{m}\rangle:=K^{T}_{m_{T}}V_{T,T-1}\dots V_{\tau+1,\tau}K^{\tau}_{m_{\tau}}V_{\tau,0}|\Psi\rangle\,.$ (8)
@@
In the sense of (12), classicality à la decoherent histories amounts to non-disturbance by third-party measurements (which is also part of macro-realism [66]). Namely, classicality emerges if the existence of measurements in $t\notin\mathbf{t}$ do not affect the outcome distribution of the ones in $t\in\mathbf{t}$. This is a reasonable way to define “classical” since only classical systems admit ideal measurements that reveal their state without changing it.
@@
$K^{t}_{m}=\frac{1}{(2\pi\Gamma^{2})^{1/4}}e^{-(\mathcal{Q}_{t}/\eta_{t}-m)^{2}/(4\Gamma^{2})}\,,m\in\mathbb{R},$ (13)
@@
Figure 2: Decoherent histories probe $\|\Delta^{\mathbf{t}}\|_{1}$ with $\mathbf{t}=\{L+\tau\}$, as function of $\tau$ (measurement starting time), for (a) $\theta=0.15\pi<\theta_{c}$ and (b) $\theta=0.3\pi>\theta_{c}$. Approximate decoherent histories emerge for coarse-grained measurements ($\Gamma=0.1$) and $\tau\gg 1$, but not for $\tau\sim 1$ or for fine-grained measurements (fg, $\Gamma\eta_{t}=0.1$). Insets (i, ii) show the pdf $p^{\mathbf{t}}$ (no third-party) and marginal distributions in non-decoherent scenarios ($T=11$). Straight lines show decay rate of $\|\Delta^{\mathbf{t}}\|_{1}$ (14), and (55) in [57]. (a) For $\theta<\theta_{c}$, $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{2\tau(x-1)}$ with no $L$ scaling. (b) For $\theta>\theta_{c}$, : $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{-T}\max(1,2\cos\theta)^{2L}$. Inset (iii): $L$-dependence is shown for two $\theta>\theta_{c}$ ($\tau=6$).
@@
Histories statistics. Although decoherent histories emerge at $t\geq\tau\gg 1$ independently of objectivity, they have distinct statistics $p_{\vec{m}}$ (9) in the two phases. In the encoding phase ($x_{1}>1/2$), the time sequence $(m_{t})$ is a Gaussian process with finite temporal correlation:
$\mathbb{E}[m_{t}]=0,\mathbb{E}[m_{t}m_{t^{\prime}}]\sim e^{-\kappa|t-t^{\prime}|}\,,\,\kappa=(x_{1}-1/2)\ln 2$ (16)
for $|t-t^{\prime}|\gg 1$. The correlation time $\kappa^{-1}$ is finite (it only diverges near the threshold $x_{1}=1/2$). To understand this intuitively, recall that the extensive observable $\mathcal{Q}_{t\gg 1}$ reveals no initial information (stored in $S$); since the expanding dynamics is self-similar, information about the system at any time $t$ becomes inaccessible at $t^{\prime}-t\gg 1$ as well.
Meanwhile, the apparatus phase ($x_{1}<1/2$) is characterized by histories with the following probabilistic law:
@@
where $\xi_{t}$ are i.i.d standard Gaussian and $\mathcal{M}$ is a random variable with the same distribution as $\mathcal{Q}_{t}/\eta_{t}$ in the long time limit (this distribution is non-Gaussian, see Fig. 2):
$\forall f\,,\,\overline{f(\mathcal{M})}=\lim_{t\to\infty}\langle V_{t,0}^{\dagger}f(\mathcal{Q}_{t}/\eta_{t})V_{t,0}\rangle.$ (18)
Namely, in every experimental run, the outcome $m_{t}$ becomes independent of $t$ up to an $O(1/\Gamma)$ noise, yet this constant value is random and varies at each run. Such behavior is akin to that of an apparatus pointer that freezes around a random value. The histories are non-ergodic: the statistical average over many runs differs from the time average over a single run. The correlation time is infinite: for any $t,t^{\prime}$, $m_{t}$ and $m_{t^{\prime}}$ are correlated. This is because $m_{t,t^{\prime}}$ is correlated with $\mathcal{Q}_{t,t^{\prime}}$ (respectively), and, as we have seen above, they are both correlated with the qubit $S$.